OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form.

OpenStudy (lgbasallote):

hint: use point slope form \[y- y_1 = m(x-x_1)\] where: m = slope x1 = x coordinate y1 = y coordinate does that help?

OpenStudy (lgbasallote):

@LordVictorius you there?

OpenStudy (anonymous):

Ya, lol

OpenStudy (anonymous):

Sorry, helping someone.

OpenStudy (lgbasallote):

so can you do it from here or do you need more help?